# COCO: Cognitive Operating System with Continuous Oversight for Multi-Agent Workflow Reliability

### v1

@@ Introduction
COCO implements a three-layer decoupled architecture that achieves intelligent monitoring and automatic correction through seamless integration with multi-agent workflows:
@@
Monitor Binding Layer: Implements loose coupling between monitoring capabilities and workflow execution through non-invasive binding mechanisms. Monitored execution nodes complete their tasks independently and then trigger monitors asynchronously, without blocking the main execution flow. The binding layer supports dynamic configuration of diverse monitoring strategies and ensures state isolation for reliable multi-instance coordination
@@
Monitoring Engine: Implements intelligent quality assessment based on large language models’ reasoning. The engine performs multi-dimensional assessment covering logical consistency, format compliance, and content completeness, transforming qualitative judgments into quantitative confidence metrics for threshold-based decisions
@@
Correction Engine: Implements the Contextual Rollback Mechanism and manages workflow state transitions. The engine preserves structured execution snapshots and integrates error diagnostic information with historical context to generate targeted correction strategies. The correction process employs the Bidirectional Reflection Protocol to guide models toward generating correct results.
@@
This architecture supports four error detection categories: logic errors, format errors, content errors, and systematic failures. COCO achieves seamless integration through annotation-based monitoring declarations that require minimal changes to existing workflow implementations.
Computational Complexity Analysis
A critical advantage of COCO is its ability to maintain quality guarantees without introducing prohibitive computational overhead. Traditional synchronous monitoring approaches introduce $O(n)$ overhead that scales linearly with workflow depth, creating performance bottlenecks that limit system throughput.
Time Complexity: COCO achieves $O(1)$ monitoring overhead relative to workflow depth through its asynchronous architecture. The monitoring process operates in parallel with workflow execution, enabling the system to maintain real-time performance while still providing comprehensive quality assurance. Key factors contributing to this efficiency include:
Asynchronous Processing: Monitor execution runs in parallel threads, avoiding execution blocking Selective Activation: Monitors are triggered only when necessary based on confidence thresholds Efficient State Management: $O(n\log{n})$ rollback optimization via dynamic programming Amortized Cost Analysis: Under typical error rates ($p<0.1$), COCO maintains amortized $O(1)$ overhead
Space Complexity: COCO’s memory requirements scale efficiently with workflow size. Checkpoint storage incurs an overhead of $O(n|S|)$, where $n$ is the workflow depth and $|S|$ is the size of the execution state. Maintaining consensus history for the Bidirectional Reflection Protocol (BRP) requires $O(w)$ memory, while the ensemble-based cross-validation mechanism introduces a fixed $O(k)$ memory overhead.
In practice, COCO’s memory usage remains within 10–15% of baseline requirements under typical parameters, making it suitable for deployment in resource-constrained environments.
Quality-Efficiency Tradeoff: COCO provides formal quality-efficiency trade-off bounds and supports dynamic monitoring intensity adjustment. System operators can configure monitoring sensitivity to balance between comprehensive error coverage and computational efficiency based on application requirements.
Task Metric
@@
GSM- Hard MMLU- Pro MBPP CommonGen- Hard AVG. Improvement Overhead
@@
Qwen3-235B-A22B 0.7210 0.7247 0.8120 0.8712 0.7822 -
Qwen3-8b 0.6808 0.5885 0.6600 0.8115 0.6852 -
Llama-3.1-8B 0.3669 0.4803 0.6020 0.7700 0.5548 -
@@
Aflow-Qwen3-8B 0.6399 0.6656 0.6500 0.8389 0.6986 -
Aflow-Llama3.1-8B 0.5648 0.4514 0.5580 0.7492 0.5809 -
@@
COCO Qwen3-8B with coco(Llama-3.1-8B) 0.6989 0.6869 0.7440 0.8450 0.7437 + 6.5%
COCO Qwen3-8B with coco(Qwen3-8B) 0.6875 0.6660 0.7660 0.8477 0.7418 + 6.2%
COCO Llama-3.1-8b with coco(Qwen3-8B) 0.6023 0.5344 0.6040 0.8028 0.6359 + 9.5%
@@
Table 1: Performance comparison of different methods across various datasets. Reference Models provide performance baselines, Multi-Agent Framework Baseline shows AFlow results, and COCO Framework demonstrates the proposed approach with different monitor configurations. Improvement overhead calculated against corresponding SOTA AFlow workflow.
Integration Efficiency: The annotation-based integration approach requires minimal codebase modifications, enabling seamless adoption across diverse workflow platforms without significant development overhead.
This analysis demonstrates that COCO achieves its quality improvements with bounded computational cost, making it a practical solution for enhancing reliability in production multi-agent systems without sacrificing performance.
System Configuration GSM-Hard Average Latency (s)
@@
Table 2: Average computing latency comparison, COCO systems are using Qwen3-8B as main model. COCO with heterogeneous configuration achieves comparable efficiency to baseline AFlow while providing quality improvements.
Experiments and Discussion
Datasets and Metrics
To thoroughly assess COCO’s error correction capabilities, we evaluate on a diverse range of logically demanding multi-agent workflow tasks covering both closed- and open-domain reasoning scenarios:
Closed-domain Tasks: These tasks require precise, objective reasoning with unambiguous correct answers, making them ideal for evaluating core error detection accuracy and correction effectiveness. GSM-Hard(Potamitis 2025) features arithmetic problems with complex multi-step calculations and large numerical values, challenging the framework’s ability to detect and correct mathematical reasoning errors while maintaining computational accuracy. MMLU-Pro(Wang 2024a) provides a comprehensive benchmark spanning diverse knowledge domains through multiple-choice questions, testing both factual recall accuracy and logical deduction correctness. Both benchmarks use accuracy as the evaluation metric, enabling precise quantification of error correction improvements.
Open-domain Tasks: These tasks are inherently creative and open-ended, requiring multi-dimensional qualitative evaluation of error correction in complex generative workflows. MBPP(Austin 2021) tasks agents with generating functionally correct code solutions, demanding proficiency in syntax error detection, logical flow correction, and semantic consistency validation. The evaluation combines code executability, test case passing rates, and algorithmic correctness, reflecting practical software development error patterns. CommonGen-Hard(Wang 2024b) challenges agents to generate coherent sentences connecting disparate concepts, testing error correction in commonsense reasoning, contextual understanding, and creative expression. Evaluation incorporates grammar accuracy, logical consistency, concept coverage, and semantic coherence, providing comprehensive assessment of generative quality under error correction.
Implementation Details
@@
AFlow Iterative Optimization Framework. We employed the AFlow framework to automatically construct and optimize multi-agent workflows through Monte Carlo Tree Search-based exploration. This process generates directed acyclic graph (DAG) structures with performance-guided agent dependencies. For fair comparative evaluation, we augmented baseline configurations with analysis nodes (the first node) to align computational complexity with COCO monitoring overhead. COCO Decorator Architecture. We implemented the monitoring system using a decorator pattern that integrates seamlessly with existing agent workflows without requiring modifications to core logic. The decorator maintains thread-local task registries, enabling clean cancellation and restart operations during error correction cycles. Importantly, the placement of COCO monitors on specific workflow nodes is manually defined based on human analysis of workflow characteristics and empirical error patterns, rather than automated dependency analysis as previously stated. Heterogeneous Multi-Agent vLLM(Zhao 2023) Parallel Processing. Our local model deployment utilizes vLLM’s asynchronous execution framework to support concurrent multi-agent processing across heterogeneous architectures:
  • Small-scale models (Qwen3-8B(Team et al. 2025), Llama-3.1-8B-Instruct-Turbo(Grattafiori et al. 2024)) run through local vLLM instances for high-throughput parallel execution.
  • Large-scale models (Qwen3-235B-A22B(Team et al. 2025)) are accessed through API interfaces due to computational constraints.
@@
For further detailed information, please refer to Appendix A.
Experimental Results
@@
Table 3: Ablation study examining individual and combined contributions of COCO’s core abilities in Bidirectional Reflection Protocol process. CoT (Chain-of-Thought) provides reasoning capabilities for the monitoring LLM. Results demonstrate synergistic effects between components.
Ablation Study Results
@@
Component Analysis. The complete COCO configuration (w/ Reflection w/ CoT) achieves optimal performance at 0.7215 average, representing a 4.6% improvement over the baseline configuration (w/o Reflection w/o CoT: 0.6900). This validates COCO’s core hypothesis: systematic error correction through the Contextual Rollback Mechanism with complete Bidirectional Reflection Protocol produces multiplicative benefits. The ablation results reveal an asymmetry between COCO’s core components—Chain-of-Thought without BRP (0.7150) significantly outperforms Bidirectional Reflection Protocol without CoT (0.6830) by 4.7%, demonstrating that enhanced reasoning capabilities provide substantial value when integrated into the CRM framework. The performance progression from baseline (0.6900) to CoT integration (0.7150) to full system (0.7215) shows that CoT integration provides substantial improvement (+3.6%), while the complete BRP implementation adds additional refinement (+0.9%) when combined with the Contextual Rollback Mechanism.
Discussion
@@
Limitations and Future Work
Technical Limitations
@@
Future Research Directions
@@ Conclusion
Beyond performance, COCO enables reliable deployment of modular AI systems in critical domains where error cascades could cause severe failures. Its annotation-based integration lowers implementation barriers, making COCO a practical solution for improving multi-agent system reliability in production.
Appendix A: Task-Specific Workflow Implementation Details
Task-Specific Implementation. We applied COCO to optimize workflows across multiple benchmark tasks, with monitor placement manually determined based on task characteristics and empirical error analysis:
GSM Hard Workflow
Baseline: The mathematical reasoning workflow begins with problem input, followed by systematic analysis to understand the problem structure and identify key variables. The system then performs custom reasoning to extract numerical values and mathematical relationships, develops computational strategies, and applies mathematical operations. Subsequently, code verification validates the mathematical logic through programmatic execution, and finally produces the complete solution with step-by-step reasoning and final numerical answer.
COCO Enhancement: We strategically placed a monitor at the custom reasoning stage, recognizing that numerical value extraction and mathematical symbol interpretation represent the most critical failure point in mathematical problem-solving. The COCO monitor continuously evaluates whether the extracted numerical content remains faithful to the original problem statement, actively detects hallucinations in mathematical interpretation, and validates the correctness of numerical formats and unit conversions before computation proceeds.
MMLU-Pro Workflow
Baseline: The knowledge-intensive reasoning workflow processes complex multi-domain questions through systematic stages. Initially, the system analyzes the question to understand domain requirements and knowledge dependencies. It then performs conflict analysis to identify potential contradictions or ambiguities in the question context. The workflow proceeds to generate an initial answer based on domain-specific reasoning, followed by ensemble generation where multiple reasoning pathways are explored in parallel. Finally, self-consistency mechanisms integrate diverse reasoning approaches to produce a confident, well-supported final answer.
COCO Enhancement: We implemented dual monitoring at two critical junctures: the initial answer generation stage and the self-consistency integration stage. The initial answer monitor operates at the critical divergence point where heterogeneous models (such as Llama and Qwen) demonstrate different reasoning approaches, enabling early detection of domain-specific knowledge gaps or logical inconsistencies. The self-consistency monitor serves as the final quality checkpoint, where reasoning errors from multiple pathways could propagate into the final output, ensuring that the integration process maintains logical coherence and factual accuracy across diverse reasoning sources.
MBPP Workflow
Baseline: The code generation workflow transforms natural language programming problems into executable solutions through a structured process. The system begins by analyzing the problem statement to understand functional requirements, input-output specifications, and algorithmic constraints. It then generates multiple candidate code solutions using diverse programming approaches and strategies. Finally, the code selector evaluates these candidates based on correctness, efficiency, and code quality metrics to choose the optimal implementation.
COCO Enhancement: We positioned a monitor at the code selection stage, recognizing this as the decisive quality gateway where algorithmic correctness and implementation excellence must be assured. The COCO monitor comprehensively evaluates candidate code solutions across multiple dimensions: logical coherence of the algorithmic approach, syntactic correctness of the implementation, semantic alignment with problem requirements, and overall problem-solving efficacy. This ensures that only robust, well-structured code solutions proceed to final output, preventing the propagation of subtle bugs, inefficient algorithms, or incomplete implementations.
CommonGen Workflow
Baseline: The creative sentence generation workflow transforms discrete concept sets into coherent, meaningful sentences through a multi-stage refinement process. The system begins by analyzing the input concepts to understand semantic relationships, thematic connections, and potential narrative structures. It then generates initial sentence candidates that attempt to incorporate all required concepts. The workflow continues with verification to check concept coverage and grammatical correctness, followed by refinement to improve coherence and readability. Finally, the selection mechanism chooses the most appropriate sentence based on fluency, concept integration, and semantic coherence criteria.
COCO Enhancement: We strategically deployed a monitor at the sentence generation stage, identifying this as the primary source of logical coherence failures based on our comprehensive error analysis. The COCO monitor performs real-time assessment of generated sentences, evaluating whether they maintain logical consistency throughout the narrative structure, ensure complete and meaningful incorporation of all required concepts, and preserve semantic coherence between disparate concept domains. This early intervention prevents the downstream propagation of logically incoherent or conceptually incomplete sentences, ensuring that subsequent refinement and selection stages operate on fundamentally sound linguistic foundations.
This strategic placement of monitoring components across all workflow architectures enables COCO to identify and correct errors at their source points, effectively preventing error propagation through downstream processing stages while maintaining system efficiency and reliability.
Appendix B: Convergence Analysis of the BRP Protocol
Convergence Proof of BRP. Under the some assumptions, the Bidirectional Reflection Protocol (BRP) terminates in finite time with probability approaching 1, i.e., it eventually produces an output that is accepted by the verifier model.
Let $\mathcal{Y}$ be the output space of candidate responses. At each interaction step $t$, an Act model samples an output $y_{t}\sim p_{t}(y\mid x_{t})$ conditioned on the context $x_{t}$. A Monitor model then evaluates this output via a binary predicate $A(y_{t},x_{t})\in\{0,1\}$, indicating whether the response is accepted. If the output fails verification (i.e., $A(y_{t},x_{t})=0$), the Monitor provides a critique $r_{t}$, which is incorporated into the subsequent input $x_{t+1}=\phi(x_{t},y_{t},r_{t})$. This process iterates until a valid output is accepted.
We aim to show that, under mild and intuitive assumptions, this iterative protocol converges with probability approaching one—that is, the interaction will terminate after finitely many steps with a high probability of acceptance.
We begin by assuming that for every possible context $x_{t}$, there exists at least one candidate output $y^{*}\in\mathcal{Y}$ such that $A(y^{*},x_{t})=1$. This assumption, denoted as H1, asserts the existence of a valid solution at every step—ensuring that the search space is not degenerate.
Next, we posit a second assumption, H2, which reflects the intuition that feedback from the Monitor is informative and corrective. More formally, if the current output $y_{t}$ is rejected, then the Monitor’s critique $r_{t}$ leads to a modified context $x_{t+1}$ such that the expected probability of generating an acceptable output in the next step strictly increases. That is,
$\mathbb{E}_{y\sim p_{t+1}(\cdot\mid x_{t+1})}[A(y,x_{t+1})]>\mathbb{E}_{y\sim p_{t}(\cdot\mid x_{t})}[A(y,x_{t})].$
Finally, we make a third assumption, H3, ensuring that the Act model has sufficient exploration capacity. Specifically, it must assign non-zero probability to all valid outputs at each step. Formally,
$\forall y^{*}\in\mathcal{Y},A(y^{*},x_{t})=1,\quad p_{t}(y^{*}\mid x_{t})>0.$
Under these three assumptions, we define the indicator variable $Z_{t}:=A(y_{t},x_{t})$ to represent whether the model’s output at step $t$ is accepted, and denote $\pi_{t}:=\mathbb{P}(Z_{t}=1)$ as the probability of acceptance at round $t$. From assumption H2, we know that if the answer is rejected, the following round’s acceptance probability improves: $\pi_{t+1}>\pi_{t}$. This yields a sequence $\{\pi_{t}\}_{t=1}^{\infty}$ that is strictly increasing whenever rejection occurs and is bounded above by 1. Thus, $\pi_{t}$ converges.
Now, since assumption H3 ensures that the Act model always retains non-zero probability of sampling a valid solution, and assumption H1 guarantees that such a solution exists at each round, it follows that once $\pi_{t}$ becomes sufficiently close to 1, the model will almost surely sample an acceptable output in finite steps. Therefore, the protocol terminates with high probability.
$\boxed{\lim_{T\to\infty}\mathbb{P}(A(y_{T},x_{T})=1)=1}$
Supporting References These assumptions, particularly the feedback-induced improvement (H2), are supported by prior work. Yuan and Xie(Yuan and Xie 2025) show that actor–critic feedback loops trained via preference optimization yield monotonic reasoning improvement. The Reflexion framework, demonstrates that verbal critique enhances multi-step correction in language agents(Shinn 2023). Additionally, RR-MP(He 2024) leverages multi-agent reflection to reduce hallucinations and stabilize reasoning, reinforcing both H2 and H3.

### v2

@@ 3.3 Overall Architecture
COCO implements a three-layer decoupled architecture that achieves intelligent monitoring and automatic correction through seamless integration with multi-agent workflows.
@@
Monitor Binding Layer: It implements loose coupling by defining a set $\tilde{\mathcal{V}}\subseteq\mathcal{V}$ through non-invasive decorators. Monitored execution nodes $A_{i}$ complete tasks $z_{i}$ independently and trigger monitors $\mathcal{M}$ asynchronously, ensuring $O(1)$ overhead without blocking execution path. This mechanism supports dynamic configuration of strategies and ensures state isolation for reliable multi-instance coordination.
@@
Monitoring Engine: The engine serves as an LLM-based analyzer $\mathcal{M}(z_{i},\Psi_{i};\theta_{mon})$ that performs multi-dimensional assessments covering logical consistency, format compliance, and content completeness. It transforms qualitative judgments into a structured diagnostic report $\mathbf{R}_{i}=\langle h_{i},c_{i},\tau_{i},\delta_{i},\gamma_{i}\rangle$, yielding quantitative confidence metrics $c_{i}$ used for threshold-based intervention.
@@
Correction Engine: This engine manages workflow state transitions by orchestrating the CRM and BRP. It preserves structured execution snapshots $\Psi_{i}$ and integrates diagnostic information $\delta_{i}$ with historical context to generate an augmented prompt $P_{i}^{*}=\mathcal{F}_{aug}(P_{i},\Delta_{i})$. The correction process utilizes the BRP to guide the model through reflective re-computation $z_{i}^{\prime}$, ensuring convergence to valid results.
@@
The architecture supports four error categories $\tau_{i}\in\{\text{logic, format, content, systematic}\}$. COCO achieves seamless integration through annotation-based monitoring declarations that require minimal changes.
@@
Task Metric GSM- Hard MMLU- Pro MBPP CommonGen- Hard AVG. Improvement Overhead
@@
Qwen3-235B-A22B 0.7210 0.7247 0.8120 0.8712 0.7822
Qwen3-8B 0.6808 0.5885 0.6600 0.8115 0.6852
Llama-3.1-8B 0.3669 0.4803 0.6020 0.7700 0.5548
@@
Aflow-Qwen3-8B 0.6399 0.6656 0.6500 0.8389 0.6986
Aflow-Llama3.1-8B 0.5648 0.4514 0.5580 0.7492 0.5809
@@
COCO Qwen3-8B with coco(Llama-3.1-8B) 0.6989 0.6869 0.7440 0.8450 0.7437 +6.50%
COCO Qwen3-8B with coco(Qwen3-8B) 0.6875 0.6660 0.7660 0.8477 0.7418 +6.20%
COCO Llama-3.1-8B with coco(Qwen3-8B) 0.6023 0.5344 0.6040 0.8028 0.6359 +9.50%
@@
Table 1: Performance benchmarking across diverse datasets. We compare COCO against standard reference models and the state-of-the-art AFlow multi-agent baseline. All improvement metrics report gains relative to the corresponding AFlow workflow.
System Configuration GSM-Hard Avg. Latency (s)
@@
Table 2: Average latency comparison with Qwen3-8B backbone. COCO’s heterogeneous configuration maintains parity with the AFlow’s efficiency while significantly enhancing task quality.
4 Experiments and Discussion
4.1 Datasets and Metrics
To assess error correction robustness, we benchmark COCO on diverse, logically demanding tasks covering both closed- and open-domain settings:
Closed-domain Tasks: These tasks require precise, objective reasoning with unambiguous correct answers, making them ideal for evaluating core error detection accuracy and correction effectiveness. GSM-HardGao et al. (2023) features arithmetic problems with complex multi-step calculations and large numerical values, challenging the framework’s ability to detect and correct mathematical reasoning errors while maintaining computational accuracy. MMLU-ProWang et al. (2024) provides a comprehensive benchmark spanning diverse knowledge domains through multiple-choice questions, testing both factual recall accuracy and logical deduction correctness. Both benchmarks use accuracy as the evaluation metric, enabling precise quantification of error correction improvements.
Open-domain Tasks: These tasks are inherently creative and open-ended, requiring multi-dimensional qualitative evaluation of error correction in complex generative workflows. MBPPChen et al. (2021) tasks agents with generating functionally correct code solutions, demanding proficiency in syntax error detection, logical flow correction, and semantic consistency validation. The evaluation combines code executability, test case passing rates, and algorithmic correctness, reflecting practical software development error patterns. CommonGen-HardMadaan et al. (2023) challenges agents to generate coherent sentences connecting disparate concepts, testing error correction in commonsense reasoning, contextual understanding, and creative expression. Evaluation incorporates grammar accuracy, logical consistency, concept coverage, and semantic coherence, providing comprehensive assessment of generative quality.
4.2 Implementation Details
@@
AFlow Iterative Optimization Framework. We employed the AFlow framework to automatically construct and optimize multi-agent workflows through Monte Carlo Tree Search-based exploration. This process generates DAG structures with performance-guided agent dependencies. For fair comparative evaluation, we augmented baseline configurations with analysis nodes (the first node) to align computational complexity with COCO monitoring overhead.
COCO Decorator Architecture. We implemented monitoring via a non-intrusive decorator pattern, enabling seamless integration without altering core workflow logic. To balance flexibility and automation, COCO employs a hybrid configuration strategy: high-risk nodes are identified either through empirical manual selection or an automated SampleAndTest protocol. This process generates an optimized overlay $\mathcal{D}$ that populates the monitoring set $\tilde{\mathcal{V}}$ for oversight.
Heterogeneous Multi-Agent vLLMKwon et al. (2023) Parallel Processing. Our local model deployment utilizes vLLM’s asynchronous framework to support concurrent multi-agent processing across heterogeneous architectures:
  • Small-scale models (Qwen3-8BYang et al. (2025), Llama-3.1-8B-Instruct-TurboGrattafiori et al. (2024)) run through local vLLM instances for high-throughput parallel execution.
  • Large-scale models (Qwen3-235B-A22BYang et al. (2025)) are accessed through API interfaces.
@@
4.3 Experimental Results
@@
Table 3: Ablation study examining individual and combined contributions of COCO’s abilities. Results demonstrate synergistic effects between Bidirectional Reflection Protocol and CoT reasoning.
4.4 Ablation Study Results
@@
Component Analysis. The complete COCO configuration (w/ Reflection w/ CoT) achieves optimal performance at 0.7215 average, representing a 4.6% improvement over the baseline configuration (w/o Reflection w/o CoT: 0.6900). This validates COCO’s core hypothesis: systematic error correction through the Contextual Rollback Mechanism with complete Bidirectional Reflection Protocol produces multiplicative benefits. The ablation results reveal an asymmetry between COCO’s core components—Chain-of-Thought without BRP (0.7150) significantly outperforms Bidirectional Reflection Protocol without CoT (0.6830) by 4.7%, demonstrating that enhanced reasoning capabilities provide substantial value when integrated into the CRM framework. The performance progression from baseline (0.6900) to CoT integration (0.7150) to full system (0.7215) shows that CoT integration provides substantial improvement (+3.6%), while the complete BRP implementation adds additional refinement (+0.9%) when combined with CRM.
4.5 Discussion
@@
5 Limitations and Future Work
5.1 Technical Limitations
@@ 5.2 Ethical Considerations
5.3 Future Research Directions
@@ 6 Conclusion
Beyond performance, COCO enables reliable deployment of modular AI systems in critical domains where error cascades could cause severe failures. Its annotation-based integration lowers implementation barriers, Furthermore, an automated configuration engine optimizes monitoring strategies via empirical sampling to minimize deployment complexity, making COCO a practical solution for improving multi-agent system reliability in production.